If the graph of y=2x^2-kx+3 touches the x-axis then what are the possible values of k?
Please keep answer in surd form.
Thanks in advance.

Answer by user_dude2008(1862)

d=b^2-4ac
0=(-k)^2-4(2)(3)
0=k^2-24
k^2=24
k=sqrt(24) or k=-sqrt(24)
k=2*sqrt(6) or k=-2*sqrt(6)

Answers: k=2*sqrt(6) or k=-2*sqrt(6)